OpenStudy (anonymous):

Find the quotient of the quantity -36x^3-6x^ + 12x over -6x. −36x3 − 6x2 − 2 −6x2 + x − 2 6x2 − x + 2 6x2 + x − 2

OpenStudy (anonymous):

6x2 + x − 2?

OpenStudy (akashdeepdeb):

See all you have to do is divide EACH term [Separated by a + or a -] by -6x \[-36x^3-6x^2 + 12x\] Now divide \[-36x^3\] by -6x^2 And then divide -6x^2 by -6x And then divide the last term by +12x by -6x You'll have your answer! :)

OpenStudy (akashdeepdeb):

YES GREAT JOB!! :D

OpenStudy (anonymous):

thanks

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

\[\frac{-36 x^3}{-6 x}-\frac{6 x^2}{-6 x}+\frac{12 x}{-6 x} \]